OpenStudy (anonymous):

Use trigonometric identities to transform one side of the equation into the other (0 < theta < pi/2) csc theta tan theta = sec theta

OpenStudy (anonymous):

I got csc theta tan theta = 1/sin theta * sin theta/cos theta = 1/cos theta = sec theta is that correct??

OpenStudy (anonymous):

\[\csc \theta \tan \theta = \sec \theta \] \[\csc \theta \tan \theta = \frac{1}{\sin \theta} * \frac{\sin \theta}{\cos \theta}\] \[= \frac{1}{\cos \theta} = \sec \theta\]

OpenStudy (anonymous):

Yes. You are correct.

OpenStudy (anonymous):

awesome thank you!! :) can you check this one really quick?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\cot \theta \sin \theta = \cos \theta\] \[\cot \theta \sin \theta = \frac{\cos \theta}{\sin \theta}*\frac{\sin \theta}{\sin \theta} = \cos \theta\]

OpenStudy (anonymous):

is this one correct?

OpenStudy (anonymous):

Yes! You are correct again! You are really smart!

OpenStudy (anonymous):

I'm trying really hard! math is not my best subject lol, thank you!! :)

OpenStudy (anonymous):

your welcome!